# Relative portfolio optimization via a value at risk based constraint

### v1

@@ 3.2. Power utility
We consider now the same problem with power utility for the agents. Note that we use the same parameter for the utility function for both agents. More precisely, we consider for $U(x)=\frac{1}{1-\gamma}x^{1-\gamma},\,\gamma\in(0,1),$ the individual portfolio problems
@@
We concentrate now on the interesting case which is determined by the parameters $\beta_{1}\beta_{2}<1$ and
@@
where $z_{1-\alpha_{2}}$ is the $(1-\alpha_{2})-$quantile of $Z_{T}$, i.e. $\mathbb{P}(Z_{T}\leq z_{1-\alpha_{2}})=1-{\alpha_{2}}$. Since $Z_{T}>0$ we necessarily have that $z_{1-\alpha_{2}}>0.$
@@
If $\beta_{1}\beta_{2}<1$ and the inequality in (3.19) holds true, the unique Nash equilibrium is given by
@@
Using Lemma 8.1 and rearranging the terms, we obtain that a Nash equilibrium is necessarily of the form
@@
We write $\lambda_{1}:=U^{\prime}(x_{0}^{1}/\varepsilon_{\gamma})$, i.e. $I(\lambda_{1})=x_{0}^{1}/\varepsilon_{\gamma}.$ It follows that $I(\lambda_{2})<\beta_{1}I(\lambda_{1})$, because otherwise the price of $X_{2}^{*}$ would be larger or equal than $\beta_{1}x_{0}^{1}$ which contradicts our assumption. This implies $\lambda_{2}>\lambda_{1}\beta_{1}^{-\gamma}$. It remains to prove that $X_{2}^{*}$ is the best response to $X_{1}^{*}$. Thus, we consider the following function
$L(X,\lambda_{2},\eta_{2}):=U(X)-\lambda_{2}Z_{T}X+\eta_{2}\mathbbm{1}\{X\geq\beta_{1}I(\lambda_{1}Z_{T})\},$
@@
$\eta_{2}:=U(I(\lambda_{2}z_{1-\alpha_{2}}))-U(\beta_{1}I(\lambda_{1}z_{1-\alpha_{2}}))+\lambda_{2}z_{1-\alpha_{2}}\Big{(}\beta_{1}I(\lambda_{1}z_{1-\alpha_{2}})-I(\lambda_{2}z_{1-\alpha_{2}})\Big{)}$
and $\lambda_{1},\lambda_{2}$ as before. Note that $\eta_{2}\geq 0$ since
$\eta_{2}\geq 0$ $\Leftrightarrow U(I(\lambda_{2}z_{1-\alpha_{2}}))-U(\beta_{1}I(\lambda_{1}z_{1-\alpha_{2}}))\geq\lambda_{2}z_{1-\alpha_{2}}\Big{(}I(\lambda_{2}z_{1-\alpha_{2}})-\beta_{1}I(\lambda_{1}z_{1-\alpha_{2}})\Big{)}$
$\Leftrightarrow\frac{U(I(\lambda_{2}z_{1-\alpha_{2}}))-U(\beta_{1}I(\lambda_{1}z_{1-\alpha_{2}}))}{I(\lambda_{2}z_{1-\alpha_{2}})-\beta_{1}I(\lambda_{1}z_{1-\alpha_{2}})}\leq\lambda_{2}z_{1-\alpha_{2}}$
@@
for some $\xi\in(I(\lambda_{2}z_{1-\alpha_{2}}),\beta_{1}I(\lambda_{1}z_{1-\alpha_{2}})).$ Because $U^{\prime}$ is non-increasing, we obtain $U^{\prime}(\xi)\leq U^{\prime}(I(\lambda_{2}z_{1-\alpha_{2}}))=\lambda_{2}z_{1-\alpha_{2}}.$ Now we can show that for any other admissible terminal wealth $X_{2}$ we obtain
@@
since $X_{2}^{*}$ maximizes $X\mapsto L(X,\lambda_{2},\eta_{2})$ pointwise (see App. 8.3) which implies that $X_{2}^{*}$ maximizes the objective of agent 2 under the VaR-based constraint. Now it only remains to discuss the existence of $\lambda_{2}>0$ such that $\mathbb{E}[Z_{T}X_{2}^{*}]=x_{0}^{2}$. We can solve
@@
Note that for power utility, in contrast to logarithmic utility, the position of the set $A$ that guarantees that the probability constraint is satisfied is not arbitrary. This is remarkable since the limit $\gamma\to 1$ results in a logarithmic utility function. Moreover, we can compare the set $\{Z_{T}>z_{1-\alpha_{2}}\}$ on which the constraint is satisfied in the power utility case to the set $\{Z_{T}\leq z_{\alpha_{2}}\}$ discussed in Remark 3.4. There we searched for the set $A_{2}$ with $\mathbb{P}(A_{2})=\alpha_{2}$ that maximizes the expected terminal wealth of agent 2 in the Nash equilibrium under logarithmic utility. On the contrary, the set $\{Z_{T}>z_{1-\alpha_{2}}\}$ found in Theorem 3.5 minimizes the expected terminal wealth discussed in Remark 3.4. Nevertheless, it does not come as a surprise that the constraint is satisfied on the set $\{Z_{T}>z_{1-\alpha_{2}}\}$ since it is ’cheaper’ to satisfy the constraint there as the terminal wealth of agent 1 is decreasing in $Z_{T}$.
@@ 4. Multi agent case
$(P_{n})\left\{\begin{array}[]{l}\mathbb{E}\ln(X_{i})\to\max\\ \mathbb{P}(X_{i}\geq\sum_{j\neq i}\beta_{j}X_{j})\geq\alpha_{i},\\ \mathbb{E}[Z_{T}X_{i}]=x_{0}^{i},\\ X_{i}\mbox{ is }\mathcal{F}-\mbox{measurable,}\end{array}\right.$
with $\alpha_{i},\beta_{i}\in[0,1]$. Further, we assume w.l.o.g. that the initial capitals of the agents are ordered by $x_{0}^{1}\geq\ldots\geq x_{0}^{n}.$ In particular, we restrict the discussion to the logarithmic utility. Moreover, we assume that for all $i$ we have $\sum_{j\neq i}\beta_{j}\leq 1.$ We are again looking for a Nash equilibrium of investment strategies. The definition for $n$ agents is as follows:
@@
4.1. Assume that $\alpha_{1}+\ldots+\alpha_{n}\leq 1$
@@
with $\lambda_{\beta}^{-i}:=\sum_{j\neq i}\beta_{j}\lambda_{j}$, $\mathbb{P}(A_{i})=\alpha_{i}$ and $A_{i}\cap A_{j}=\emptyset$ for $i\neq j.$ Moreover, we have that
@@
$(P_{n})\left\{\begin{array}[]{l}\mathbb{E}\ln(X_{i})\to\max\\ X_{i}\geq\mathbbm{1}_{A_{i}}\sum_{j\neq i}\beta_{j}X_{j},\\ \mathbb{E}[Z_{T}X_{i}]=x_{0}^{i},\\ X_{i}\mbox{ is }\mathcal{F}-\mbox{measurable,}\\ A_{i}\in\mathcal{F},\mathbb{P}(A_{i})=\alpha_{i}.\end{array}\right.$
@@
where $X_{\beta}^{-i}:=\sum_{j\neq i}\beta_{j}X_{j}$. The maximum construction and the self-financing constraint yield $X_{i}\geq\lambda_{i}/Z_{T}$ and $\lambda_{i}\leq x_{0}^{i}.$ Thus, we obtain that the minimal value of $Z_{T}X_{\beta}^{-i}$ is attained on the set $(\cup_{j\neq i}A_{j})^{c}.$ Taking Lemma 8.2 into account, in order to maximize $\mathbb{E}\ln(X_{i})$ we have to choose $A_{i}$ such that $A_{i}\cap A_{j}=\emptyset$ for $i\neq j.$ Due to the assumption $\alpha_{1}+\ldots+\alpha_{n}\leq 1$, this is possible and the sets $A_{i}$ are disjoint in a Nash equilibrium. In particular we obtain
@@
$X_{2}$ $=$ $\mathbbm{1}_{A_{2}}\max\bigg{\{}\beta_{1}X_{1}+\sum_{j\geq 3}\beta_{j}X_{j},\frac{\lambda_{2}}{Z_{T}}\bigg{\}}+\mathbbm{1}_{A_{2}^{c}}\frac{\lambda_{2}}{Z_{T}}.$
@@
$X_{2}=\mathbbm{1}_{A_{2}}\max\bigg{\{}\frac{\beta_{1}\lambda_{1}}{Z_{T}}+\sum_{j\geq 3}\beta_{j}X_{j},\frac{\lambda_{2}}{Z_{T}}\bigg{\}}+\mathbbm{1}_{A_{2}^{c}}\frac{\lambda_{2}}{Z_{T}}.$ (4.4)
@@
Solving this equation for $\lambda_{i}$ yields (4.3). In order to have a viable solution, we must have $\lambda_{i}\geq 0$ which yields the inequalities. ∎
@@
    Since $\lambda_{i}\leq x_{0}^{i}$ and $\sum_{j\neq i}\beta_{j}\leq 1$ we always have that $x_{0}^{1}\geq\bar{\lambda}_{-1}$ and thus $X_{1}^{*}=\frac{x_{0}^{1}}{Z_{T}}$. Hence, the richest agent will never be influenced by the probability constraint.
4.2. Assume that $\alpha_{1}+\ldots+\alpha_{n}>1$
@@
$(PD_{i})\left\{\begin{array}[]{l}\sum_{k=1}^{m}\ln(\lambda_{ki})\to\max\\ \lambda_{ki}\geq M_{ki}\sum_{j\neq i}\beta_{j}\lambda_{kj},\;\mbox{ for all }k,\\ \sum_{k=1}^{m}\lambda_{ki}=mx_{0}^{i},\\ \sum_{k=1}^{m}M_{ki}=\ell_{i},\\ \lambda_{ki}>0,M_{ki}\in\{0,1\},\;\mbox{ for all }k.\end{array}\right.$
@@
$X_{i}=\frac{1}{Z_{T}}\Big{(}\sum_{k\in S}\max\big{\{}\sum_{j\neq i}\beta_{j}\lambda_{kj},\lambda\big{\}}\mathbbm{1}_{B_{k}}+\sum_{k\in S^{c}}\lambda\mathbbm{1}_{B_{k}}\Big{)}$
where $S$ is a set of $\ell_{i}$ indices where $\sum_{j\neq i}\beta_{j}\lambda_{kj}$ is smallest over all $k.$ Thus, we can write $X_{i}$ as in (4.6) with $\lambda_{ki}$ either being $\sum_{j\neq i}\beta_{j}\lambda_{kj}$ or $\lambda.$ The self-financing condition can be stated as $\mathbb{E}[Z_{T}X_{i}]=1/m\sum_{k=1}^{m}\lambda_{ki}=x_{0}^{i}.$ The probability constraint is satisfied when for $\ell_{i}$ indices $\lambda_{ki}\geq\sum_{j\neq i}\beta_{j}\lambda_{kj}.$ Since $\ln(x)$ is Schur-concave the optimal indices are automatically chosen by the optimization problem. ∎
5. Beating a stochastic benchmark
We would also like to mention that the general setting we discussed here may be specified to a classical benchmark problem where we have only one agent who tries to beat a (stochastic) benchmark with at least a certain probability. We explain this with a simple specific example: Suppose we have a Black Scholes market with one riskless bond with zero interest and one stock with price process
$\mathrm{d}S_{t}=$ $S_{t}\Big{(}\mu\,\mathrm{d}t+\sigma\,\mathrm{d}W_{t}\Big{)},$
where $S_{0}=1$, $\sigma>0$ and $\mu\in\mathbb{R}$ are given parameters with $\sigma^{2}>\mu$. W.l.o.g. we assume that $S_{0}=1.$ Thus we have
$S_{t}=$ $\exp\Big{(}\Big{(}\mu-\frac{1}{2}\sigma^{2}\Big{)}t+\sigma W_{t}\Big{)},$
$Z_{t}=$ $\exp\Big{(}-\frac{1}{2}\frac{\mu^{2}}{\sigma^{2}}t-\frac{\mu}{\sigma}W_{t}\Big{)}$
where $(Z_{t})$ is again the state price density. We consider the problem
$(P_{B})\left\{\begin{array}[]{l}\mathbb{E}\ln(X)\to\max\\ \mathbb{P}(X\geq\beta S_{T})\geq\alpha,\\ \mathbb{E}[Z_{T}X]=x_{0},\\ X\mbox{ is }\mathcal{F}-\mbox{measurable.}\\ \end{array}\right.$
Thus, we maximize the logarithmic utility of terminal wealth under the constraint that we beat at least with probability $\alpha$ the benchmark $\beta S_{T}.$ The initial wealth is $x_{0}.$ We can use Lemma 8.2 to derive an explicit solution. First, the optimal set where the constraint will be satisfied is given by
$M_{\lambda_{\alpha}}:=\Big{\{}\beta S_{T}\leq\frac{\lambda_{\alpha}}{Z_{T}}\Big{\}}\mbox{ where }\lambda_{\alpha}\mbox{ is such that }\mathbb{P}(M_{\lambda_{\alpha}})=\alpha.$ (5.1)
In our application this set can be characterized further. Note that
$S_{T}Z_{T}=(Z_{T})^{1-\frac{\sigma^{2}}{\mu}}\exp\Big{(}\frac{1}{2}(\mu-\sigma^{2})T\Big{)}.$ (5.2)
Thus, when we define
$\kappa(\lambda):=\Big{(}\frac{\lambda}{\beta}\Big{)}^{\frac{\mu}{\mu-\sigma^{2}}}\exp\Big{(}-\frac{1}{2}\mu T\Big{)},$ (5.3)
we can express $M_{\lambda_{\alpha}}$ as $M_{\lambda_{\alpha}}=\{Z_{T}\geq\kappa(\lambda_{\alpha})\}.$ Thus, in total we obtain with Lemma 8.2 that the optimal terminal wealth is $X^{*}=x_{0}/Z_{T}$ if $x_{0}\geq\lambda_{\alpha}$. In the case $x_{0}<\lambda_{\alpha}$, we obtain
$X^{*}=\left\{\begin{array}[]{cl}\lambda/Z_{T},&\mbox{ if }Z_{T}\leq\kappa(\lambda_{\alpha})\mbox{ or }Z_{T}>\kappa(\lambda)\\ \beta S_{T},&\mbox{ if }Z_{T}\in(\kappa(\lambda_{\alpha}),\kappa(\lambda)]\end{array}\right.$
where $\lambda>\lambda_{\alpha}$ is such that $\mathbb{E}[X^{*}Z_{T}]=x_{0}.$ This problem is related to a number of similar questions which have been treated in the literature before such as maximizing the probability of beating a benchmark or minimizing shortfall (see e.g. Browne,, 1999; Korn and Lindberg,, 2014; Föllmer and Schied,, 2016).

### v2

@@ 3.3. Power utility
We consider now the same problem with power utility for the agents. Note that we use the same parameter for the utility function for both agents. More precisely, we consider for $U(x)=\frac{1}{1-\gamma}x^{1-\gamma},\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\gamma>0,\gamma\neq 1},$ the individual portfolio problems
@@
We concentrate now on the interesting case which is determined by the parameters $\beta_{1}\beta_{2}<1$ and for $\gamma\in(0,1)$ by
@@
and for $\gamma>1$ by
[…]
where $z_{\alpha}$ is the $\alpha-$quantile of $Z_{T}$, i.e. $\mathbb{P}(Z_{T}\leq z_{\alpha})=\alpha$. Since $Z_{T}>0$ we necessarily have that $z_{\alpha}>0.$
@@
Let $\beta_{1}\beta_{2}<1$.
  • a)
    If $\gamma\in(0,1)$ and the inequality in (3.20) holds true, the unique Nash equilibrium is given by
@@
  • b)
    If $\gamma>1$ and the inequality in (3.21) holds true, the unique Nash equilibrium is given by
    $X_{1}^{*}=\frac{x_{0}^{1}}{\varepsilon_{\gamma}}Z_{T}^{-1/\gamma},\mbox{ and }X_{2}^{*}=\left\{\begin{array}[]{cr}(\lambda_{2}Z_{T})^{-1/\gamma},&Z_{T}>z_{\alpha_{2}},\\ \beta_{1}\frac{x_{0}^{1}}{\varepsilon_{\gamma}}Z_{T}^{-1/\gamma},&Z_{T}\leq z_{\alpha_{2}},\end{array}\right.$ (3.23)
    where $\lambda_{2}>0$ is such that $\mathbb{E}[Z_{T}X_{2}^{*}]=x_{0}^{2}$.
@@
We restrict the proof to case a). Case b) is similar. Using Lemma 7.1 and rearranging the terms, we obtain that a Nash equilibrium is necessarily of the form
@@
We write $\kappa:=U^{\prime}(x_{0}^{1}/\varepsilon_{\gamma})$, i.e. $I(\kappa)=x_{0}^{1}/\varepsilon_{\gamma}.$ It follows that $I(\lambda_{2})<\beta_{1}I(\kappa)$, because otherwise the price of $X_{2}^{*}$ would be larger or equal than $\beta_{1}x_{0}^{1}$ which contradicts our assumption. This implies $\lambda_{2}>\kappa\beta_{1}^{-\gamma}$. It remains to prove that $X_{2}^{*}$ is the best response to $X_{1}^{*}$. Thus, we consider the following function
$L(X,\lambda_{2},\eta_{2}):=U(X)-\lambda_{2}Z_{T}X+\eta_{2}\mathbbm{1}\{X\geq\beta_{1}I(\kappa Z_{T})\},$
@@
$\eta_{2}:=U(I(\lambda_{2}z_{1-\alpha_{2}}))-U(\beta_{1}I(\kappa z_{1-\alpha_{2}}))+\lambda_{2}z_{1-\alpha_{2}}\Big{(}\beta_{1}I(\kappa z_{1-\alpha_{2}})-I(\lambda_{2}z_{1-\alpha_{2}})\Big{)}$
and $\kappa,\lambda_{2}$ as before. Note that $\eta_{2}\geq 0$ since
$\eta_{2}\geq 0$ $\Leftrightarrow U(I(\lambda_{2}z_{1-\alpha_{2}}))-U(\beta_{1}I(\kappa z_{1-\alpha_{2}}))\geq\lambda_{2}z_{1-\alpha_{2}}\Big{(}I(\lambda_{2}z_{1-\alpha_{2}})-\beta_{1}I(\kappa z_{1-\alpha_{2}})\Big{)}$
$\Leftrightarrow\frac{U(I(\lambda_{2}z_{1-\alpha_{2}}))-U(\beta_{1}I(\kappa z_{1-\alpha_{2}}))}{I(\lambda_{2}z_{1-\alpha_{2}})-\beta_{1}I(\kappa z_{1-\alpha_{2}})}\leq\lambda_{2}z_{1-\alpha_{2}}$
@@
for some $\xi\in(I(\lambda_{2}z_{1-\alpha_{2}}),\beta_{1}I(\kappa z_{1-\alpha_{2}})).$ Because $U^{\prime}$ is non-increasing, we obtain $U^{\prime}(\xi)\leq U^{\prime}(I(\lambda_{2}z_{1-\alpha_{2}}))=\lambda_{2}z_{1-\alpha_{2}}.$ Now we can show that for any other admissible terminal wealth $X_{2}$ we obtain
@@
since $X_{2}^{*}$ maximizes $X\mapsto L(X,\lambda_{2},\eta_{2})$ pointwise (see App. 7.3) which implies that $X_{2}^{*}$ maximizes the objective of agent 2 under the VaR-based constraint. Now it only remains to discuss the existence of $\lambda_{2}>0$ such that $\mathbb{E}[Z_{T}X_{2}^{*}]=x_{0}^{2}$. We can solve
@@
  1. a)
    Note that for power utility, in contrast to logarithmic utility, the position of the set $A$ that guarantees that the probability constraint is satisfied is not arbitrary. This is remarkable since the limit $\gamma\to 1$ results in a logarithmic utility function. Moreover, it turns out that for $\gamma>1$ the constraint is satisfied on the set $\{Z_{T}\leq z_{\alpha_{2}}\}$ whereas for $\gamma\in(0,1)$, the constraint is satisfied on the set $\{Z_{T}>z_{1-\alpha_{2}}\}$. In both cases the constraint is satisfied on the side where it is cheaper to achieve. This is the main reason for the non-continuous behavior of the optimal terminal wealth as a function of $\gamma.$ It cannot be explained by the risk aversion. The relative risk aversion of the power utility is given by $-x\frac{U^{\prime\prime}(x)}{U^{\prime}(x)}=\gamma$ and thus risk aversion is increasing in $\gamma$. When we consider as an example again the Black Scholes market, it holds that large $Z_{T}$ correspond to small stock prices and vice versa. This means that the less risk-averse investor (less risk averse than log-utility) tries to outperform the other agent in the event of small stock prices, whereas the more risk-averse investor (more risk averse than log-utility) tries to outperform the other agent in the event of large stock prices.
  2. b)
    It is possible to derive the replicating strategies for the Nash equilibrium from Theorem 3.6 using Theorem E.1 from Jin and Yu Zhou, (2008). However, the resulting expressions become even more complex than the ones obtained for logarithmic utility which is why we refrain from presenting them here.
@@ 4. Multi agent case
$(P_{n})\left\{\begin{array}[]{l}\mathbb{E}\ln(X_{i})\to\max\\ \mathbb{P}(X_{i}\geq\sum_{j\neq i}\beta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ij}}X_{j})\geq\alpha_{i},\\ \mathbb{E}[Z_{T}X_{i}]=x_{0}^{i},\\ X_{i}\mbox{ is }\mathcal{F}-\mbox{measurable,}\end{array}\right.$
with $\alpha_{i},\beta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ij}}\in[0,1]$. Further, we assume w.l.o.g. that the initial capitals of the agents are ordered by $x_{0}^{1}\geq\ldots\geq x_{0}^{n}.$ In particular, we restrict the discussion to the logarithmic utility. Moreover, we assume that for all $i$ we have $\sum_{j\neq i}\beta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ij}}\leq 1.$ We are again looking for a Nash equilibrium of investment strategies. The definition for $n$ agents is as follows:
@@
4.1. Assume that the sum of the alphas is less or equal to 1
@@
with $\lambda_{\beta}^{-i}:=\sum_{j\neq i}\beta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ij}}\lambda_{j}$, $\mathbb{P}(A_{i})=\alpha_{i}$ and $A_{i}\cap A_{j}=\emptyset$ for $i\neq j.$ Moreover, we have that
@@
$(P_{n})\left\{\begin{array}[]{l}\mathbb{E}\ln(X_{i})\to\max\\ X_{i}\geq\mathbbm{1}_{A_{i}}\sum_{j\neq i}\beta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ij}}X_{j},\\ \mathbb{E}[Z_{T}X_{i}]=x_{0}^{i},\\ X_{i}\mbox{ is }\mathcal{F}-\mbox{measurable,}\\ A_{i}\in\mathcal{F},\mathbb{P}(A_{i})=\alpha_{i}.\end{array}\right.$
@@
where $X_{\beta}^{-i}:=\sum_{j\neq i}\beta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ij}}X_{j}$. The maximum construction and the self-financing constraint yield $X_{i}\geq\lambda_{i}/Z_{T}$ and $\lambda_{i}\leq x_{0}^{i}.$ Thus, we obtain that the minimal value of $Z_{T}X_{\beta}^{-i}$ is attained on the set $(\cup_{j\neq i}A_{j})^{c}.$ Taking Lemma 7.2 into account, in order to maximize $\mathbb{E}\ln(X_{i})$ we have to choose $A_{i}$ such that $A_{i}\cap A_{j}=\emptyset$ for $i\neq j.$ Due to the assumption $\alpha_{1}+\ldots+\alpha_{n}\leq 1$, this is possible and the sets $A_{i}$ are disjoint in a Nash equilibrium. In particular we obtain
@@
$X_{2}$ $=$ $\mathbbm{1}_{A_{2}}\max\bigg{\{}\beta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}21}}X_{1}+\sum_{j\geq 3}\beta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2j}}X_{j},\frac{\lambda_{2}}{Z_{T}}\bigg{\}}+\mathbbm{1}_{A_{2}^{c}}\frac{\lambda_{2}}{Z_{T}}.$
@@
$X_{2}=\mathbbm{1}_{A_{2}}\max\bigg{\{}\frac{\beta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}21}}\lambda_{1}}{Z_{T}}+\sum_{j\geq 3}\beta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2j}}X_{j},\frac{\lambda_{2}}{Z_{T}}\bigg{\}}+\mathbbm{1}_{A_{2}^{c}}\frac{\lambda_{2}}{Z_{T}}.$ (4.4)
@@
Solving this equation for $\lambda_{i}$ yields (4.3). In order to have a viable solution, we must have $\lambda_{i}\geq 0$ which yields the inequalities.
@@
    Since $\lambda_{i}\leq x_{0}^{i}$ and $\sum_{j\neq i}\beta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ij}}\leq 1$ we always have that $x_{0}^{1}\geq\bar{\lambda}_{-1}$ and thus $X_{1}^{*}=\frac{x_{0}^{1}}{Z_{T}}$. Hence, the richest agent will never be influenced by the probability constraint.
4.2. Assume that the sum of the alphas is larger than 1
@@
$(PD_{i})\left\{\begin{array}[]{l}\sum_{k=1}^{m}\ln(\lambda_{ki})\to\max\\ \lambda_{ki}\geq M_{ki}\sum_{j\neq i}\beta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ij}}\lambda_{kj},\;\mbox{ for all }k,\\ \sum_{k=1}^{m}\lambda_{ki}=mx_{0}^{i},\\ \sum_{k=1}^{m}M_{ki}=\ell_{i},\\ \lambda_{ki}>0,M_{ki}\in\{0,1\},\;\mbox{ for all }k.\end{array}\right.$
@@
$X_{i}=\frac{1}{Z_{T}}\Big{(}\sum_{k\in S}\max\big{\{}\sum_{j\neq i}\beta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ij}}\lambda_{kj},\lambda\big{\}}\mathbbm{1}_{B_{k}}+\sum_{k\in S^{c}}\lambda\mathbbm{1}_{B_{k}}\Big{)}$
where $S$ is a set of $\ell_{i}$ indices where $\sum_{j\neq i}\beta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ij}}\lambda_{kj}$ is smallest over all $k.$ Thus, we can write $X_{i}$ as in (4.6) with $\lambda_{ki}$ either being $\sum_{j\neq i}\beta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ij}}\lambda_{kj}$ or $\lambda.$ The self-financing condition can be stated as $\mathbb{E}[Z_{T}X_{i}]=1/m\sum_{k=1}^{m}\lambda_{ki}=x_{0}^{i}.$ The probability constraint is satisfied when for $\ell_{i}$ indices $\lambda_{ki}\geq\sum_{j\neq i}\beta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ij}}\lambda_{kj}.$ Since $\ln(x)$ is Schur-concave the optimal indices are automatically chosen by the optimization problem. ∎